OpenStudy (anonymous):

Simplify: (3x^3 y^4)^2

OpenStudy (anonymous):

9x^6y^8.

OpenStudy (anonymous):

9x^5y^6 6x^6y^8 9x^6y^8 none of the above

OpenStudy (anonymous):

... 9x^6y^8.

OpenStudy (anonymous):

\[\Large \begin{array}{l} {\left( {3{x^3}{y^4}} \right)^2}\\ = \left( {3{x^3}{y^4}} \right)\left( {3{x^3}{y^4}} \right)\\ = 3{x^3} \cdot 3{x^3} \cdot {y^4} \cdot {y^4}\\ = 9{x^{3 + 3}}{y^{4 + 4}}\\ = 9{x^6}{y^8} \end{array}\]